# Near-real-time flood inundation monitoring by Bayesian analysis for change point problems for Sentinel-1 time series

### v2

@@ Abstract
Near real-time flood monitoring is crucial for disaster response, yet existing methods face significant limitations in training data requirements and cloud cover interference. Here we present a novel approach using Bayesian analysis for change point problems (BCP) applied to Sentinel-1 SAR time series data, which automatically detects temporal discontinuities in backscatter patterns to distinguish flood inundation from permanent water bodies without requiring training data or ancillary information. We validate our method using the UrbanSARFloods benchmark dataset across three diverse geographical contexts (Weihui, China; Jubba, Somalia; and NovaKakhovka, Ukraine). Our BCP approach achieves F1 scores ranging from 0.41 to 0.76 (IoU: 0.25-0.61), significantly outperforming both OTSU thresholding (F1: 0.03-0.12, IoU: 0.02-0.08) and Siamese convolutional neural network approaches (F1: 0.08-0.34, IoU: 0.05-0.24). Further analysis reveals exceptional performance in open areas with F1 scores of 0.47-0.81 (IoU: 0.31-0.68) and high recall (0.36-0.84), contrasted with substantially lower performance in urban areas (F1: 0.00-0.01, IoU: 0.00-0.01), indicating a common challenge across current flood detection methods in urban environments. The proposed method’s ability to process raw SAR data directly with minimal preprocessing enables integration into operational early warning systems for rapid flood mapping, particularly in agricultural and open landscapes where it demonstrates the strongest performance.
keywords:
Synthetic Aperture Radar , Flood monitoring , Earth observation
journal: Remote Sensing of Environment
[…]
[not,tt,ns]organization=Graduate school of Science & Engineering, Saitama University, addressline=255 Shimo-Okubo, city=Sakura ward, Saitama, postcode=338-8570, state=Saitama, country=Japan
{graphicalabstract}[Uncaptioned image]
{highlights}
Novel Bayesian analysis for change point problems detects floods in Sentinel-1 time series
Method requires no training data or ancillary info to distinguish flood from permanent water
Automated processing of raw SAR data enables near real-time flood monitoring applications
@@ 1 Introduction
Flooding is one of the most frequent and devastating natural disasters globally [Alfieri2017-ma, Najibi2018-ta], resulting in substantial loss of life, damage to infrastructure, and severe economic consequences [Kundzewicz2014-ge, Jongman2021-ea, Tellman2021-hv, Jonkman2024-hi, Rentschler2021-dc]. The economic impact and human toll of flooding necessitate effective monitoring systems for early detection and rapid response [Oddo2019-or, Rentschler2020-pm]. Near real-time flood inundation mapping, enabled by advances in remote sensing, has emerged as a critical tool for effective disaster response [Oddo2019-or, Katiyar2021-fy, Shen2019-aj, Tay2020-cp]. These systems provide situational awareness to disaster responders, guide evacuation planning and search and rescue operations, and enable rapid damage assessments [Browder2021-io]. However, the widespread adoption of near real-time flood mapping remains limited due to challenges in remote sensing data analytics, creating an urgent need for more efficient and accessible monitoring approaches.
@@
Satellite remote sensing offers significant advantages for flood monitoring, including wide spatial coverage, high temporal resolution, and the ability to observe remote or inaccessible areas. However, optical remote sensing techniques face substantial limitations due to their inability to penetrate cloud cover, which is often present during flood events. Synthetic Aperture Radar (SAR) addresses these limitations by operating in the microwave portion of the electromagnetic spectrum, enabling all-weather, day-night imaging capability. The launch of the Sentinel-1 mission by the European Space Agency (ESA) has greatly enhanced the availability of SAR data, providing global coverage with a revisit time of 6-12 days and free data access. With high spatial resolution (5-20 m) and wide swath width (250 km), Sentinel-1 is well-suited for mapping flood extents at regional to continental scales [Schlaffer2015-uv].
Despite SAR’s advantages for flood monitoring, significant challenges persist in developing fully automated, efficient, and accurate flood detection methods. Many existing approaches rely on supervised classification techniques that require extensive training data [Shastry2023-wl, Wu2023-kc, Ghosh2024-ub], which may not be readily available for all flood scenarios. Supervised classification methods require large amounts of labeled data to train algorithms, creating significant barriers to operational flood monitoring. The transferability of trained models to new flood scenarios or geographical regions is often limited [Zhao2024-lv], requiring additional data collection and model retraining. Furthermore, reliance on ancillary data (land cover maps, digital elevation models, hydrological information) can restrict the applicability and timeliness of flood mapping approaches, particularly in rapidly changing environments or regions with limited data infrastructure. Most existing time-series approaches for SAR-based flood monitoring rely either on simple change detection between pre- and post-flood images or on conventional statistical measures that don’t fully leverage the temporal information content in SAR data. Traditional change detection methods struggle to account for seasonal variations in backscatter or to distinguish between different types of change (e.g., flooding versus agricultural harvesting). While some Bayesian approaches have been applied in other remote sensing applications [Barry1993-qp, Erdman2008-kv], they have not been effectively adapted to exploit the unique characteristics of SAR time series for flood detection, particularly the need to identify abrupt changes against background variability.
@@
Thus, this study aims to overcome these limitations by developing an efficient, automated solution for near real-time flood monitoring using Sentinel-1 SAR time series data. We propose a pixel-wise Bayesian statistical model for SAR-based flood mapping that operates without the need for training data. Unlike existing approaches, our method leverages a Bayesian analysis of Change Point problem (BCP) as proposed by Barry and Hartigan [Barry1993-qp] to identify flood-induced changes in the mean of our SAR time series data for the first time. Our workflow distinguishes between permanent water bodies and temporary flood inundations without requiring ancillary data by leveraging the temporal change characteristics in the probability domain rather than the raw backscatter domain. We use raw Sentinel-1 Ground Range Detected (GRD) time series as input without requiring training data, pre-trained models, or extensive manual intervention. The method achieves robust performance across diverse geographical contexts, making it particularly suitable for operational flood monitoring applications.
@@ 2 Related works
2.1 Flood monitoring by Optical remote sensing data
Optical satellite platforms such as Sentinel-2 and PlanetScope offer high spatial resolution and multispectral capabilities for flood mapping [Portales-Julia2023-pz, Wieland2019-je]. However, cloud cover significantly impedes continuous monitoring during critical initial flood phases. Portalés-Julià et al. (2023) found approximately 50% of pixels obscured by clouds during the first post-flood satellite revisit [Portales-Julia2023-pz]. Recent deep learning approaches have attempted to estimate water bodies beneath clouds [Li2021-fq, Shastry2023-wl]. These models leverage spatial context from adjacent cloud-free pixels and temporal dependencies between acquisitions, achieving over 90% accuracy in water classification under 50-80% cloud coverage [Portales-Julia2023-pz]. Despite these advances, challenges remain regarding training data requirements and geographical transferability.
2.2 Flood monitoring by remote sensing SAR data
SAR presents a compelling alternative due to its all-weather imaging capability [Landuyt2019-cz, Amitrano2024-av, Shen2019-aj]. SAR discriminates between water and non-water surfaces based on distinctive backscattering properties, with smooth water surfaces typically appearing dark due to specular reflection.
@@ 2.2 Flood monitoring by remote sensing SAR data
Interferometric SAR (InSAR) utilizes phase coherence information between acquisitions to detect inundation patterns, demonstrating effectiveness in complex urban environments [Chini2019-me]. However, the practical application of InSAR for timely flood monitoring is hindered by significant challenges. Normally, the extensive processing needed to derive these parameters often prevents real-time flood monitoring, limiting the method’s operational utility in urgent disaster response scenarios.
Deep learning methods have demonstrated exceptional capabilities for SAR-based flood detection [Saleh2024-vd, Bereczky2022-qy], with recent studies reporting classification accuracies exceeding 98% in certain contexts [Ghosh2024-ub]. Specialized datasets like Sen1Flood11 [Bonafilia2020-sk] and Kuro-Siwo [Bountos2023-jg] have facilitated model training across diverse geographical contexts.
@@
Nevertheless, significant challenges persist: (1) extensive labeled training data requirements, (2) limited model transferability across geographical contexts [Wu2023-kc], (3) difficulty distinguishing between permanent water bodies and temporary inundation [Notti2018-vs], and (4) complex backscatter interpretation in urban or vegetated areas [Amitrano2024-av].
2.3 Time Series Analysis for Flood Detection
@@
Recent research has increasingly leveraged the temporal dimension of satellite data for improved flood detection [Schlaffer2015-uv]. Time series approaches exploit the sequential nature of acquisitions to identify anomalous patterns indicative of flood events, particularly effective with Sentinel-1’s consistent 6-12 day revisit cycle [Oddo2019-or]. Schlaffer et al. (2015) proposed a harmonic analysis approach using multi-temporal SAR data to detect floods based on deviations from expected seasonal patterns [Schlaffer2015-uv]. Similarly, Tsyganskaya et al. (2018) and Tsyganskaya et al. (2019) employed time series analysis for wetland change detection and flooded vegetation mapping using Sentinel-1 data [Tsyganskaya2018-ts, Tsyganskaya2019-nv]. These methods enable more robust discrimination between permanent water bodies and temporary inundation.
@@
Conventional Bayesian methods offer advantages for hydrological applications by incorporating prior knowledge and quantifying uncertainty [Barry1993-qp, Erdman2008-kv]. D’Addabbo et al. (2016) developed a Bayesian network model fusing SAR and ancillary data [D-Addabbo2016-yi], and Giordano et al. (2023) applied Bayesian statistical modeling to multi-temporal Sentinel-1 stacks for high-resolution flood monitoring [Giordano2023-ys]. However, existing Bayesian approaches primarily focus on retrospective analysis rather than near real-time detection. Most methods lack specific adaptations for the characteristic SAR backscatter patterns during flood events—sharp decreases during inundation followed by gradual increases during recession, limiting their effectiveness for timely flood monitoring.
@@ 2.4 Research Gap and Our Contribution
Previous studies faces challenges in SAR-based flood monitoring that our approach addresses: (1) extensive training data requirements limiting applicability in data-scarce regions, (2) domain adaptation difficulties across flood scenarios, and (3) challenges distinguishing permanent water bodies from flood inundation. Our approach with Bayesian analysis for change point problems (BCP) operates without training data, processes Sentinel-1 time series directly, and leverages temporal patterns to differentiate water classes without pre-existing water masks. This approach fills a critical gap in flood monitoring studies, offering a complementary method particularly valuable when timely detection is essential for emergency response.
@@ 3 Methods
The workflow of our approach consists of three main components: (1) Reference data and Sentinel-1 time series data preparation, (2) BCP analysis, and (3) post-processing with spatial filtering. The workflow is designed to process raw SAR data directly and provide automated flood extent mapping without requiring training data or extensive manual intervention.
@@ 3.1.1 UrbanSARFloods
As reference data, we utilized the UrbanSARFloods dataset, a comprehensive benchmark for flood detection using Sentinel-1 SAR data [Zhao2024-lv]. This dataset encompasses 8,879 non-overlapping image chips of 512×512 pixels at 20-m spatial resolution, covering an area of 807,500 km2 across 18 flood events on 5 continents. Every pixel in this dataset is classified into three distinct categories: non-flooded areas (0), flooded open areas (1), and flooded urban areas (2). UrbanSARFloods is systematically divided into training (4,501 chips), validation (1,970 chips), and testing (2,408 chips) sets, which are independent of flood events in training and validation data. This dataset ensures its quality by a two-tier annotation process, which combines semi-automatic labeling through conventional remote sensing techniques with manual verification using high-resolution optical data from PlanetScope and UAV imagery. While the BCP analysis does not require any training/validation set, we used them for the comparative analysis described in Section 4.
To evaluate the geographical robustness of our method, we conducted extensive testing across three diverse test sites found in UrbanSARFloods: Weihui (China), Jebba (Somalia), and NovaKakhovka (Ukraine). Weihui represents an urban environment with complex infrastructure and dense building patterns. In such settings, SAR backscatter is influenced by double-bounce reflection mechanisms between water surfaces and vertical structures, creating complex scattering patterns that can confound conventional flood detection approaches. The regular grid pattern of roads and buildings creates distinctive backscatter signatures that require careful interpretation. Jebba exemplifies a rural, agricultural landscape with varying topography and seasonal vegetation patterns. In these areas, the temporal variability of vegetation growth can mask or mimic flood signatures, particularly in partially submerged vegetation where volume scattering mechanisms dominate. The predominantly open terrain presents different challenges compared to urban environments, with larger homogeneous regions but greater seasonal variability. NovaKakhovka represents a mixed environment with both urban elements and natural landscapes, including significant riverine features. This combination of built environment and natural terrain presents a particularly challenging test case, as algorithms must correctly distinguish permanent water bodies (river) from newly flooded areas while accounting for different backscatter behaviors across varied land cover types.
@@ 3.1.2 Sentinel-1 GRD
Sentinel-1 GRD offers VV and VH polarization modes that measure radar backscatter amplitude using C-band SAR instruments. The data undergoes multi-looking processing and is projected onto the WGS84 ellipsoid, converting radar measurements into ground coordinates for geographic analysis. This multi-looking process reduces noise while preserving backscatter amplitude, though it removes phase information. For each flood event, we collected time series data spanning one year before the event using Google Earth Engine, to establish a robust baseline for change detection. We didn’t take into account of further processing such as speckle filtering and radiometric terrain normalization to ensure the sensitivity of time-series variability of backscatter intensity to changes in surface properties. To account for differences in viewing geometry and illumination conditions over time, we filtered the timeseries data in the platform type (A or B) and the orbit direction (ascending or descending) that influences backscatter response due to varying incidence angles and shadow effects based on the observation at the flood event (Table 1). While Sentinel-1 GRD data is 10-m spatial resolution, we aggregated into 20-m spatial resolution to match with UrbanSARFloods in this experiment.
Table 1: Tested flood events and observation condition during the flood event.

### v3

# Automated flood detection from Sentinel-1 GRD time series using Bayesian analysis for change point problems

@@ Abstract
Current Synthetic Aperture Radar (SAR)-based flood detection methods face critical limitations that hinder operational deployment. Supervised learning approaches require extensive labeled training data, exhibit poor geographical transferability, and may fail to adapt to new regions without additional training examples. Existing approaches do not fully exploit the rich temporal information available in SAR time series, instead relying on simple change detection between pre- and post-flood images or supplementary datasets that often introduce error propagation. These limitations prevent effective automated flood monitoring in data-scarce regions where disaster response is most needed. To address these limitations, we develop a novel training-free approach by adapting Bayesian analysis for change point problems, specifically for automated flood detection from Sentinel-1 Ground Range Detected time series data. Our method statistically models the temporal behavior of SAR backscatter intensity over a one-year baseline period, then computes the posterior probability of change points at flood observation dates. This approach eliminates supervised learning dependencies by using Bayesian inference to identify when backscatter deviations exceed expected normal variations, leveraging inherent statistical properties of time series data. Validation across three diverse geographical contexts using the UrbanSARFloods benchmark dataset demonstrates superior performance compared to conventional thresholding and deep learning approaches, achieving F1 scores up to 0.76. This enables immediate deployment to any region with SAR coverage, providing critical advantages for disaster response.
Keywords Synthetic Aperture Radar $\cdot$ Flood monitoring $\cdot$ Earth observation
@@ 2 Introduction
Flooding is one of the most frequent and devastating natural disasters globally [1, 2], resulting in substantial loss of life, damage to infrastructure, and severe economic consequences [3, 4, 5, 6, 7]. The economic impact and human toll of flooding necessitate effective monitoring systems for early detection and rapid response [8, 9]. While satellite remote sensing has emerged as a critical tool for flood mapping [8, 10, 11, 12], providing situational awareness to disaster responders and enabling damage assessments [13], the widespread adoption of operational flood mapping remains limited due to fundamental methodological challenges in remote sensing data analytics.
@@
Synthetic Aperture Radar (SAR) offers significant advantages for flood monitoring through its ability to operate in the microwave portion of the electromagnetic spectrum, enabling all-weather, day-night imaging capability regardless of cloud cover. The Sentinel-1 mission by the European Space Agency (ESA) has greatly enhanced the availability of SAR data, providing global coverage with a revisit time of 6-12 days, free data access, high spatial resolution (5-20 m), and wide swath width (250 km), making it well-suited for mapping flood extents at regional to continental scales [14]. While Sentinel-1 data is available in both Single Look Complex (SLC) and Ground Range Detected (GRD) formats, most operational applications utilize GRD products due to their reduced preprocessing requirements and accessibility for operational monitoring systems.
Despite these advantages, current SAR-based flood mapping approaches face significant methodological limitations. Many existing techniques rely on supervised classification methods that require extensive training data [15, 16, 17], creating substantial barriers to operational flood monitoring in data-scarce regions. The transferability of trained models to new flood scenarios or geographical contexts is often limited [18, 19], necessitating additional data collection and model retraining for each new application area. A further critical limitation is the failure to fully exploit the rich temporal information available in SAR time series. Current approaches frequently depend on supplementary datasets such as optical satellite imagery [20, 21], digital elevation models [22], or hydrological information, creating multi-source integration challenges that increase computational overhead, introduce error propagation, and compromise the temporal consistency needed for real-time flood monitoring.
To address these limitations, some studies have focused on developing analytical frameworks that fully exploit time-series approaches for Sentinel-1 GRD-based flood monitoring. However, these methods remain methodologically constrained, relying primarily on simple change detection between pre- and post-flood images or on conventional statistical measures that inadequately capture the temporal dynamics of backscatter behavior. Traditional change detection methods struggle to distinguish flood-induced backscatter changes from natural seasonal variations or other land surface modifications (e.g., agricultural practices, vegetation phenology), often resulting in high false alarm rates. Furthermore, these methods typically require manual threshold selection and extensive parameter tuning, hindering their automation potential. While Bayesian time series analysis methods have shown promise for detecting abrupt changes [23, 24], they have not been effectively adapted to exploit the unique temporal characteristics of SAR backscatter for flood detection. This represents a significant research gap, as SAR time series inherently contain the temporal context necessary to distinguish flood events from background variability without requiring external data sources—offering the potential for fully automated flood monitoring systems.
@@
This study aims to overcome these fundamental methodological challenges by developing an automated solution for flood detection using Sentinel-1 GRD time series data. We propose a pixel-wise Bayesian statistical model for SAR-based flood mapping that operates without the need for training data or pre-existing water masks. Unlike existing approaches, our method leverages a Bayesian analysis of Change Point problems (BCP) as proposed by Barry and Hartigan [23] to identify flood-induced changes in Sentinel-1 GRD time series data. Our workflow distinguishes between permanent water bodies and temporary flood inundations without requiring ancillary data by leveraging the temporal change characteristics in the probability domain rather than the backscatter domain. By working directly with readily available GRD products, our approach eliminates the complex preprocessing requirements associated with SLC data while still achieving robust flood detection performance. By addressing key methodological limitations in current flood detection approaches, our BCP-based technique offers an advancement in the field of SAR-based flood monitoring.
@@ 3 Related works
3.1 Flood monitoring by remote sensing data
@@
While optical satellite platforms such as Sentinel-2 and PlanetScope offer high spatial resolution and multispectral capabilities for flood mapping [25, 26, 27], cloud cover significantly impedes continuous monitoring during critical initial flood phases [27]. SAR presents a compelling alternative due to its all-weather imaging capability [28, 29, 30, 11]. SAR discriminates between water and non-water surfaces based on distinctive backscattering properties, with smooth water surfaces typically appearing dark due to specular reflection. Interferometric SAR (InSAR) utilizes phase coherence information between acquisitions to detect inundation patterns [31], demonstrating effectiveness in complex urban environments [32, 33]. However, the practical application of InSAR for timely flood monitoring is hindered by significant challenges. Normally, the extensive processing needed to derive these parameters often prevents real-time flood monitoring, limiting the method’s operational utility in urgent disaster response scenarios.
@@
Deep learning methods have demonstrated exceptional capabilities for SAR-based flood detection [34, 35], with recent studies reporting classification accuracies exceeding 98% in certain contexts [17]. Specialized datasets like Sen1Floods11 [36], UrbanSARFloods [18], and Kuro Siwo [37] have facilitated model training across diverse geographical contexts. Nevertheless, significant challenges persist from SAR time series: (1) extensive labeled training data requirements, (2) limited model transferability across geographical contexts [16], (3) difficulty distinguishing between permanent water bodies and temporary inundation [38, 21], and (4) complex backscatter interpretation in urban or vegetated areas [30].
3.2 SAR Time Series for Flood Detection
@@
Recent research has increasingly leveraged the temporal dimension of SAR data for improved flood detection [14]. Time series approaches exploit the sequential nature of acquisitions to identify anomalous patterns indicative of flood events, particularly effective with Sentinel-1’s consistent 6-12 day revisit cycle [8]. Schlaffer et al. (2015) proposed a harmonic analysis approach using multi-temporal SAR data to detect floods based on deviations from expected seasonal patterns [14]. Similarly, Tsyganskaya et al. (2018) [39] and Tsyganskaya et al. (2019) [40] employed time series analysis for wetland change detection and flooded vegetation mapping using Sentinel-1 data. These methods enable more robust discrimination between permanent water bodies and temporary inundation.
@@
Conventional Bayesian methods offer advantages for hydrological applications by incorporating prior knowledge and quantifying uncertainty [23, 24]. D’Addabbo et al. (2016) [41] developed a Bayesian network model fusing SAR and ancillary data, and Giordano et al. (2023) [42] applied Bayesian statistical modeling to multi-temporal Sentinel-1 stacks for high-resolution flood monitoring. However, existing Bayesian approaches primarily focus on retrospective analysis rather than up-to-date detection of the latest observation. Most methods lack specific adaptations for the characteristic SAR backscatter patterns during flood events. These patterns typically include sharp decreases during inundation followed by gradual increases during recession, which limits their effectiveness for timely flood monitoring.
@@ 3.3 Research Gap and Our Contribution
Previous studies face challenges in SAR-based flood monitoring that our approach addresses: (1) extensive training data requirements limiting applicability in flood detection, (2) domain adaptation difficulties across flood scenarios, and (3) challenges distinguishing permanent water bodies from flood inundation. Our approach with BCP operates without training data, processes Sentinel-1 GRD time series directly, and leverages temporal patterns to differentiate water classes without pre-existing water masks. This approach fills a critical gap in flood monitoring studies, offering a complementary method particularly valuable when timely detection is essential for emergency response.
@@ 4 Methods
The workflow of our approach consists of three main components: (1) reference data and Sentinel-1 GRD time series data preparation, (2) BCP analysis, and (3) post-processing with spatial filtering (Figure 1). The workflow is designed to process GRD data directly and provide automated flood extent mapping without requiring training data or extensive manual intervention.
Figure 1: Methodological flowchart comparing three flood detection methods: Bayesian analysis of change point problems (BCP), Otsu’s thresholding, and Siamese CNN. Blue boxes indicate external datasets, including input Sentinel-1 data and the UrbanSARfloods evaluation dataset.
@@ 4.1.1 UrbanSARFloods
As reference data, we utilized the UrbanSARFloods dataset [18], a comprehensive benchmark for flood detection using Sentinel-1 SAR data. This dataset encompasses 8,879 non-overlapping image chips of 512×512 pixels at 20-m spatial resolution, covering an area of 807,500 km2 across 18 flood events on 5 continents. Every pixel in this dataset is classified into three distinct categories: non-flooded areas, flooded open areas, and flooded urban areas. UrbanSARFloods is systematically divided into training (4,501 chips), validation (1,970 chips), and testing (2,408 chips) data. The testing data are independent of flood events in the training and validation data. This dataset ensures its quality by a two-tier annotation process, which combines semi-automatic labeling through conventional remote sensing techniques with manual verification using high-resolution optical data from PlanetScope and UAV imagery [18]. While the BCP analysis does not require any training/validation data, we used them for the comparative analysis described in Section 5.
To evaluate the geographical robustness of our method, we conducted extensive testing across three diverse test sites found in UrbanSARFloods: Weihui (China), Jebba (Somalia), and Nova Kakhovka (Ukraine) (Figure 2). Weihui represents an urban environment with complex infrastructure and dense building patterns. In such settings, SAR backscatter is influenced by double-bounce reflection mechanisms between water surfaces and vertical structures, creating complex scattering patterns that can confound conventional flood detection approaches. The regular grid pattern of roads and buildings creates distinctive backscatter signatures that require careful interpretation. Jebba exemplifies a rural, agricultural landscape with varying topography and seasonal vegetation patterns. In these areas, the temporal variability of vegetation growth can mask or mimic flood signatures, particularly in partially submerged vegetation where volume scattering mechanisms dominate. The predominantly open terrain presents different challenges compared to urban environments, with larger homogeneous regions but greater seasonal variability. Nova Kakhovka represents a mixed environment with both urban elements and natural landscapes, including significant riverine features. This combination of built environment and natural terrain presents a particularly challenging test case, as algorithms must correctly distinguish permanent water bodies (river) from newly flooded areas while accounting for different backscatter behaviors across varied land cover types.
Figure 2: Overview map of tested flood events for (A) Weihui, China, (B) Jebba, Somalia, and (C) Nova Kakhovka, Ukraine.
@@ 4.1.2 Sentinel-1 GRD
Sentinel-1 GRD offers VV and VH polarization modes that measure radar backscatter amplitude using C-band SAR instruments. The data undergoes multi-looking processing and is projected onto the WGS84 ellipsoid, converting radar measurements into ground coordinates for geographic analysis. This multi-looking process reduces noise while preserving backscatter amplitude, though it removes phase information. For each flood event, we collected time series data spanning one year before the event using Google Earth Engine, to establish a robust baseline for change detection. We didn’t take into account of further processing such as speckle filtering and radiometric terrain normalization to ensure the sensitivity of time-series variability of backscatter intensity to changes in surface properties before applying the BCP. To account for differences in viewing geometry and illumination conditions over time, we filtered the timeseries data in the platform type (A or B) and the orbit direction (ascending or descending) that influences backscatter response due to varying incidence angles and shadow effects based on the observation at the flood event (Table 1).
While Sentinel-1 GRD data is originally available at 10-m spatial resolution, we deliberately aggregated it to 20-m resolution to match the UrbanSARFloods dataset. This methodological decision was critical because aggregating results from 10-m to 20-m after processing would potentially mask pixel-level misclassifications at the original resolution. By performing the aggregation before analysis, we ensured a more accurate and fair comparison between our predictions and the reference data.
Table 1: Tested flood events and observation conditions during the flood event.